OpenStudy (anonymous):

Multiply each pair of factors. Type the product in the space provided. (6 + 3i)(6 − 3i) = (4 − 5i)(4 + 5i) = (−3 + 8i)(−3 − 8i) =

OpenStudy (anonymous):

\[\large \color{green}{(a+bi)(a-bi) = a^2 + b^2}\]

OpenStudy (anonymous):

Try to solve them with this formula.. I do it for first part: In the first part, you can compare both and you get : \(a =6\) and \(b=3\), so answer will be : \(a^2 + b^2 = 6^2 + 3^2 = \color{blue}{45}..\)